OpenStudy (anonymous):

the altitude of a triangle is 6cm greater than its base. if its area is 108 square cm, find its base

OpenStudy (wolf1728):

Area = (1/2) base * height 108 = .5 * base * (base+6) 108 = .5*base^2 + 6*base 108 = base^2/2 + 3*base base^2/2 + 3*base -108 = 0 Using quadratic formula: base =12 altitude = 18

OpenStudy (anonymous):

im sorry but I don't know what you did can you explain

OpenStudy (anonymous):

@wolf1728